OpenStudy (anonymous):

for a spherical object what is its mass in terms of the density of the material, its diameter and any necessary constants?

OpenStudy (lgbasallote):

well volume of a sphere is \[\frac 43 \pi r^3\]

OpenStudy (lgbasallote):

and then density = mass/volume

OpenStudy (unklerhaukus):

\[\rho=\frac mV\] \[m=V\rho\]

OpenStudy (lgbasallote):

so you have \[m = \frac 43\rho \pi r^3\]

OpenStudy (lgbasallote):

where \(\rho\) = density

OpenStudy (anonymous):

what about its diameter

OpenStudy (lgbasallote):

change r to d/2

OpenStudy (unklerhaukus):

and \(r\) is the radius (half the diameter) \(r=\frac d2\)

OpenStudy (lgbasallote):

radius = diameter/2 so just change r to d/2

OpenStudy (anonymous):

ok so thats the final solution?

OpenStudy (lgbasallote):

replace r with d/2....

OpenStudy (anonymous):

ok cool and can i give you guys both metals or does just one of you get it?

OpenStudy (anonymous):

hey so its (4/3)p(pi)(d/2)^3

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

is that right?

OpenStudy (unklerhaukus):

you can simplify the numbers a bit

OpenStudy (anonymous):

unkle...it won't let me give you a metal now

OpenStudy (anonymous):

why cant i give it to both of you

OpenStudy (unklerhaukus):

i dont mind about that

OpenStudy (unklerhaukus):

\[m=\frac43\rho\pi\left(\frac d2\right)^3\] use \[\left(\frac{a}{b}\right)^n=\frac {a^n} {b^n}\]

OpenStudy (anonymous):

so it simplifies to 4p(pi)(d/2)?

OpenStudy (anonymous):

i crossed out the 3's

OpenStudy (lgbasallote):

what he meant was distribute the exponent

OpenStudy (unklerhaukus):

you can not cancel an index with denominator like that

OpenStudy (anonymous):

oh so whats it turn to?

OpenStudy (unklerhaukus):

\[\left(\frac{a}{b}\right)^n=\frac {a^n} {b^n}\] \[\left(\frac{d}{2}\right)^3=\frac {d^3} {2^3}\]

OpenStudy (anonymous):

thanks :)

OpenStudy (unklerhaukus):

can you simplify the numbers in the equation now?

OpenStudy (anonymous):

yes (4/3)p(pi)(d^3/8) ?

OpenStudy (lgbasallote):

you can simplify it further

OpenStudy (lgbasallote):

you can multiply the denominators

OpenStudy (anonymous):

so 4p(pi)(d/8) ? ?

OpenStudy (lgbasallote):

what happened?

OpenStudy (anonymous):

can you simplify it?

OpenStudy (lgbasallote):

you canceled the denominaotr and teh exponent???

OpenStudy (anonymous):

it can't be simplified?

OpenStudy (lgbasallote):

it can...you just simplified it wrong....multiply the denominators

OpenStudy (lgbasallote):

don't touch anything else

OpenStudy (anonymous):

so it (4/24)p(pi)(d^3) ?

OpenStudy (unklerhaukus):

thats better!

OpenStudy (lgbasallote):

yes. now express 4/24 in lowest terms

OpenStudy (anonymous):

ok so 1/6 p(pi)(d^3) ? thats final!!!!

OpenStudy (lgbasallote):

right!!

OpenStudy (lgbasallote):

your turn @UnkleRhaukus

OpenStudy (unklerhaukus):

\[\color{red}\checkmark\]

OpenStudy (unklerhaukus):

so the constants necessary to express the volume in terms of diameter are ; the natural number; 6 the density of the material rho; \(\rho\) the geometrical constant pi; \(\pi\) the diameter of the sphere \(d\) and the dimensional constant \(3\)